# Data-adaptive structural change-point detection via isolation

### v1

@@ 4.1 Computational complexity
We will now explain why DAIS needs to check at most $\lceil\frac{T}{\lambda_{T}}\rceil+2N+1$ intervals before the algorithm is terminated, where $N$ is the total number of change-points of the signal, $T$ is its length and $\lambda_{T}$ the chosen expansion parameter. If there are no change-points, the algorithm will check at most $\lceil\frac{T}{\lambda_{T}}\rceil+1$ left- and right-expanding intervals around the location of the largest difference, as explained in Section 2. In the case that the signal has change-points, since every time detection occurs the already used observations are discarded and the algorithm is restarted on two disjoint subintervals, the total number of expansions required increases by 2 (1 for each of the newly created subintervals). Therefore, the maximum number of iterations depends on the number of change-points and DAIS needs to check at most $K^{\prime}=\lceil\frac{T}{\lambda_{T}}\rceil+2N+1$ intervals. Since we take $\lambda_{T}<\delta_{T}$, then $K^{\prime}>\lceil T/\delta_{T}\rceil+2N+1$. Combining this with the definition of $\delta_{T}$, which implies that $N\leq T/\delta_{T}$, it can be concluded that the lower bound of $K^{\prime}$ is $\mathcal{O}(T/\delta_{T})$. As explained by [2], the lower bound for the maximum number of intervals that need to be checked in ID is $\mathcal{O}(T/\delta_{T})$ and in WBS and NOT it is $\mathcal{O}(T^{2}/\delta_{T}^{2})$ up to a logarithmic factor, which means that DAIS is faster than the latter two algorithms and is of the same order as ID.
Now, focusing on the algorithms ID and DAIS, we note that the ideas behind the two methods are similar, since both first isolate the change-point and then detect it. However, the data-adaptive nature of DAIS leads to computational advancements. Using again the definition of $\delta_{T}$ and the equation $\delta_{T}\geq 2n\lambda_{T}$, for which the explanation is in Section 2.3, we get
$N\leq\frac{T}{2n\lambda_{T}}<\frac{T}{2\lambda_{T}}-1.$
So, $K^{\prime}<2T/\lambda_{T}$ and since ID checks at most $2\lceil\frac{T}{\lambda_{T}}\rceil$ intervals, it holds that the maximum number of intervals that are required for DAIS to check a signal is always less than the maximum number of intervals that ID requires. For example, when the signal has no change-points, using the same expansion parameter $\lambda_{T}$, DAIS checks almost half as many intervals as ID.
4.2 Parameter selection and Variants
@@
For the choice of the positive threshold constant $C$, we ran an extensive simulation study using various signal structures and Gaussian noise. The best behavior occurred for $C=1.7$ in the case of piecewise-constant signals and $C=2.1$ in the case of piecewise-linear signals. These are the constants that were used for the simulations in Section 5. [2] proves in Corollary 1 that, in the case of piecewise-constant signals, as $T\rightarrow\infty$, the threshold can be taken to be at most $\sqrt{3\log T}$, meaning that $C\leq\sqrt{3}\approx 1.73$; our choice does not violate this result.
@@
Variants: In practice, we use a modification to the algorithm as described in Section 2, in order to make it more accurate. DAIS restarts from the detected change-point instead of the start- and end-points of the interval where the detection occurred. This leads to the improvement of the accuracy of the algorithm without affecting the speed.
@@ 5 Simulations
This section compares the performance of DAIS with state-of-the-art competitors. We only focus on algorithms that estimate both the number and the locations of the change-points in the given univariate signal in an offline manner. The competitors used in the simulation study can be found in Table 2, along with the R library from which each function was obtained. For ID and WBS we report the results for both the information criterion and the thresholding (for $C=1.56$ for ID and $C=\sqrt{2}$ for WBS) stopping rules. The notation is ID_ic, WBS_ic and ID_th, WBS_th, respectively. In order to allow the competitors achieve their best performance, in some cases inputs of the functions as developed in R were adjusted. The competitors ID, WBS2, PELT, MOSUM and MARS were applied using the default values. The value for the maximum number of change-points to be detected was altered in each signal appropriately for WBS and NOT. For MSCP, the value of the minimal window considered is set to be less than the smallest distance between consecutive change-points. The estimation of the standard deviation for CPOP is set to be done using the MAD method and TF is employed based on the implementation found in https://github.com/hadley/l1tf. For TS, the algorithm for a continuous piecewise-linear signal with independent observations is used. Code on the simulations, as well as on how to implement DAIS, can be found in https://github.com/Sophia-Loizidou/DAIS.
@@
WBS [18] wbs
Piecewise-constant WBS2 [20] breakfast
NOT [5] not
@@
We measure the difference between the number of change-points detected and the true number of change-points ($\hat{N}-N$). As a measure of accuracy of the location of the detected change-points, we give Monte-Carlo estimates of the mean squared error,
@@
where $T$ is the total length of the sequence. For piecewise-constant signals, $\hat{f_{t}}$ is the ordinary least square approximation of $f_{t}$ between two successive change-points while for piecewise-linear signals it is the splines fit. We also report the scaled Hausdorff distance, defined as
@@
where $n_{s}=\max_{j=1,\ldots,N}\delta_{j}$ is the length of the largest segment between successive change-points. The Hausdorff distance, $d_{H}$, is given for all signals except for signals that do not contain any change-points; in such cases, $d_{H}$ is uninformative. Finally, the average time taken for each method to find the change-points is reported. The signals used, can be found in Appendix A and some further simulations results are in Appendix C.
Table 3 concerns a signal with no change-points. Most algorithms have very good results with negligible differences in the accuracy in the number of change-points detected, WBS_th with the threshold constant $C=\sqrt{2}$ seems to overestimate. Similar conclusions hold for Table 4, in which the big length of the signal (S2) forces some algorithms, especially MSCP, to be extremely slow. Moreover, comparing DAIS with ID_th which use the same value for the expansion parameter $\lambda_{T}$, it is evident that DAIS requires a smaller number of expansions and thus is faster.
@@
The signal (S3) (Table 5) is a particularly difficult structure for most algorithms because it includes pairs of consecutive change-points that shift the data sequence to opposite directions, leading to masking the existence of the change-points. Detection is more difficult when the change-points are close to each other and their isolation is unlikely or occurs only in a limited number of intervals. The best performing methods are DAIS and MOSUM, having $9\%$ improvement in the number of times the correct number of change-points is detected compared to the other algorithms. The reason why DAIS performs very well is that it will most likely start the expansion around the change-points and so it has an advantage in detecting them, as the power of the contrast function $C_{s,e}^{b}(\boldsymbol{X})$ is maximized when $b$ is at the midpoint of $[s,e]$. In Tables 6 and 7, the majority of the algorithms perform well; however, WBS_th and MSCP seem to over- and under-estimate, respectively, the number of the change-points. In Table 8, some methods struggle to detect the change-points, which are really close to each other. Based on the simulations, we conclude that overall, DAIS performs very well in various different piecewise-constant signal scenarios, involving the number of change-points, the distance between them, as well as the magnitudes of the changes, when most competitors struggle to accurately detect the changes in some of the cases tested.
@@
Proceeding now with the piecewise-linear framework, the simulations results in Table 9, show that most algorithms perform well, with the exception of MARS and TF. Tables 10, 11, 12 show that in the case of piecewise-linear signals, DAIS performs as well as the algorithms ID_ic and CPOP, and better than other competitors available in the literature. The average computational time for DAIS is very low, and most of the times, significantly lower than the best-performing competitors.
@@
Table 3: Distribution of $\hat{N}-N$ of 100 simulated data sequences of the signal (S1). The average MSE and computational time are also given.
@@
Method 0 1 2 $\geq 3$ MSE Time (s)
DAIS 99 0 1 0 $1.87\times 10^{-4}$ 0.196
ID_th 84 8 7 1 $5.26\times 10^{-4}$ 0.268
ID_ic 99 1 0 0 $1.68\times 10^{-4}$ 0.063
WBS_th 22 16 31 31 $31.7\times 10^{-4}$ 0.079
WBS_ic 99 1 0 0 $1.68\times 10^{-4}$ 0.079
WBS2 91 5 3 1 $2.75\times 10^{-4}$ 3.250
PELT 100 0 0 0 $1.39\times 10^{-4}$ 0.002
NOT 100 0 0 0 $1.39\times 10^{-4}$ 0.062
MOSUM 94 2 4 0 $2.96\times 10^{-4}$ 0.028
MSCP 100 0 0 0 $1.39\times 10^{-4}$ 83.250
Table 4: Distribution of $\hat{N}-N$ of 100 simulated data sequences of the signal (S2). The average MSE, $d_{H}$ and computational time are also given.
$\hat{N}-N$
Method $-1$ 0 1 2 $\geq 3$ MSE $d_{H}$ Time (s)
DAIS 0 99 0 1 0 $4.33\times 10^{-4}$ $43.10\times 10^{-4}$ 0.322
ID_th 0 88 1 10 1 $7.77\times 10^{-4}$ $35.80\times 10^{-4}$ 0.416
ID_ic 0 100 0 0 0 $4.25\times 10^{-4}$ $1.09\times 10^{-4}$ 0.110
WBS_th 0 43 14 24 19 $16.00\times 10^{-4}$ $186.00\times 10^{-4}$ 0.140
WBS_ic 0 100 0 0 0 $4.19\times 10^{-4}$ $1.06\times 10^{-4}$ 0.140
WBS2 0 92 3 4 1 $5.46\times 10^{-4}$ $17.60\times 10^{-4}$ 5.950
PELT 0 100 0 0 0 $4.19\times 10^{-4}$ $1.06\times 10^{-4}$ 0.004
NOT 0 100 0 0 0 $4.19\times 10^{-4}$ $1.06\times 10^{-4}$ 0.103
MOSUM 0 92 2 6 0 $5.49\times 10^{-4}$ $203.60\times 10^{-4}$ 0.049
MSCP 6 93 1 0 0 $341.35\times 10^{-4}$ $334.74\times 10^{-4}$ 276.294
Table 5: Distribution of $\hat{N}-N$ of 100 simulated data sequences of the signal (S3). The average MSE, $d_{H}$ and computational time are also given.
$\hat{N}-N$
Method $-2$ $-1$ 0 1 2 $\geq 3$ MSE $d_{H}$ Time (s)
DAIS 15 2 79 2 2 0 0.015 0.096 0.004
ID_th 2 1 72 9 13 3 0.015 0.079 0.004
ID_ic 28 0 70 2 0 0 0.016 0.154 0.002
WBS_th 0 0 18 17 19 46 0.028 0.259 0.022
WBS_ic 29 0 70 1 0 0 0.016 0.159 0.022
WBS2 4 3 72 10 6 5 0.015 0.090 0.471
PELT 80 0 20 0 0 0 0.027 0.413 0.001
NOT 30 0 68 2 0 0 0.016 0.164 0.023
MOSUM 10 1 80 5 4 0 0.013 0.083 0.011
MSCP 84 15 1 0 0 0 0.031 0.442 157.422
Table 6: Distribution of $\hat{N}-N$ of 100 simulated data sequences of the signal (S4). The average MSE, $d_{H}$ and computational time are also given.
$\hat{N}-N$
Method $\leq-3$ $-2$ $-1$ 0 1 2 $\geq 3$ MSE $d_{H}$ Time (s)
@@
MSCP 100 0 0 0 0 0 0 0.954 0.182 4.830
Table 7: Distribution of $\hat{N}-N$ of 100 simulated data sequences of the signals (S5). The average MSE, $d_{H}$ and computational time are also given.
$\hat{N}-N$
Method $\leq-3$ $-2$ $-1$ 0 1 2 $\geq 3$ MSE $d_{H}$ Time (s)
@@
MSCP 47 35 16 2 0 0 0 10.730 0.108 19.422
Table 8: Distribution of $\hat{N}-N$ of 100 simulated data sequences of the signal (S6). The average MSE, $d_{H}$ and computational time are also given.
@@
MSCP 0 100 0 3.944 0.093 92.196
Table 9: Distribution of $\hat{N}-N$ of 100 simulated data sequences of the signal (S7). The average MSE and computational time are also given.
@@
Method 0 1 2 $\geq 3$ MSE Time (s)
DAIS 100 0 0 0 $1.95\times 10^{-3}$ 0.011
ID_th 83 10 4 3 $2.94\times 10^{-3}$ 0.030
ID_ic 100 0 0 0 $1.95\times 10^{-3}$ 0.007
CPOP 100 0 0 0 $1.95\times 10^{-3}$ 2.830
NOT 100 0 0 0 $1.95\times 10^{-3}$ 0.062
MARS 0 100 0 0 $5.22\times 10^{-3}$ 0.003
TF 7 44 37 12 $2.87\times 10^{-3}$ 0.419
TS 100 0 0 0 $1.95\times 10^{-3}$ 0.502
Table 10: Distribution of $\hat{N}-N$ of 100 simulated data sequences of the signal (S8). The average MSE, $d_{H}$ and computational time are also given.
@@
Method $-1$ 0 1 $2$ $\geq 3$ MSE $d_{H}$ Time (s)
DAIS 0 99 1 0 0 0.032 0.099 0.009
ID_th 0 76 20 4 0 0.039 0.153 0.01
ID_ic 0 96 4 0 0 0.030 0.113 0.007
CPOP 0 99 1 0 0 0.013 0.056 5.390
NOT 0 99 1 0 0 0.016 0.063 0.185
MARS 2 9 42 39 8 0.034 0.200 0.006
TF 0 0 0 0 100 0.018 0.451 0.736
TS 1 95 4 0 0 0.095 0.185 0.854
Table 11: Distribution of $\hat{N}-N$ of 100 simulated data sequences of the signal (S9). The average MSE, $d_{H}$ and computational time are also given.

### v2

@@ 4.1 Computational complexity
We will now explain why DAIS needs to check at most $\lceil\frac{T}{\lambda_{T}}\rceil+2N+1$ intervals before the algorithm is terminated, where, as explained before, $N$ is the total number of change-points of the signal, $T$ is the length of the data sequence and $\lambda_{T}$ is the chosen expansion parameter. If there are no change-points, the algorithm will check at most $\lceil\frac{T}{\lambda_{T}}\rceil+1$ left- and right-expanding intervals around the location of the largest difference, as explained in Section 2. In the case that the signal has change-points, since at every time that detection occurs the already used observations are discarded and the algorithm is restarted on two disjoint subintervals, then the maximum total number of expansions required increases by two (one for each of the newly created subintervals). Therefore, the maximum number of iterations depends on the number of change-points and DAIS needs to check at most $K^{\prime}=\lceil\frac{T}{\lambda_{T}}\rceil+2N+1$ intervals. Since we take $\lambda_{T}<\delta_{T}$, then $K^{\prime}>\lceil T/\delta_{T}\rceil+2N+1$. Combining this with the definition of $\delta_{T}$, which implies that $N\leq T/\delta_{T}$, it can be concluded that the lower bound of $K^{\prime}$ is $\mathcal{O}(T/\delta_{T})$. As explained by [2], the lower bound for the maximum number of intervals that need to be checked in ID is $\mathcal{O}(T/\delta_{T})$ and in WBS and NOT it is $\mathcal{O}(T^{2}/\delta_{T}^{2})$ up to a logarithmic factor, which means that DAIS is faster than the latter two algorithms and is of the same order as ID.
Now, focusing more on the algorithms ID and DAIS, we need to highlight that the data-adaptive nature of DAIS leads to computational advancements. More specifically, due to the methodology of ID, when the data sequence under consideration has a small number of true change-points, the number of intervals that need to be checked will be close to the maximum possible for ID, which is $2\lceil\frac{T}{\lambda_{T}}\rceil$. In contrast, DAIS needs to check only $K^{\prime}$ intervals, which will be much smaller compared to ID. For example, when the signal has no change-points, using the same expansion parameter $\lambda_{T}$, DAIS checks almost half as many intervals as ID. This can be seen in the simulation results of signals (S11) and (S12) in Table 8 in Appendix B, which have 0 and 1 change-points, respectively, when compared with ID_th, which uses the same $\lambda_{T}$ as DAIS. When many change-points are present, using again the definition of $\delta_{T}$ and the inequality $\delta_{T}\geq 2n\lambda_{T}$, for which the explanation is in Section 2.3, we get $N\leq T/(2n\lambda_{T})<T/(2\lambda_{T})-1.$ So, $K^{\prime}$ will always be smaller than the maximum number of intervals that are checked by ID. However, it needs to be noted that the data-adaptivity comes at an additional, but small, computational cost, since, at each step of the algorithm, the location of the largest difference, in absolute value, needs to be re-calculated.
4.2 Parameter selection
@@
For the choice of the positive threshold constant $C$, we ran an extensive simulation study using various signal structures and Gaussian noise. The best behaviour occurred for $C=1.7$ in the case of piecewise-constant signals and $C=2.1$ in the case of piecewise-linear signals. These are the constants that were used for the simulations in Section 5. [2] proves in Corollary 1 that, in the case of piecewise-constant signals, as $T\rightarrow\infty$, the threshold can be taken to be at most $\sqrt{3\log T}$, meaning that $C\leq\sqrt{3}\approx 1.73$; our choice does not violate this result.
@@ 5 Simulations
This section compares the performance of DAIS with state-of-the-art competitors. We only focus on algorithms that estimate both the number and the locations of the change-points in the given univariate signal in an offline manner. The competitors used in the simulation study can be found in Table 2, along with the relevant R library. For ID, WBS and SeedBS we report the results for both the information criterion and the thresholding (for the default values $C=1.56$ for ID and $C=\sqrt{2}$ for WBS and SeedBS) stopping rules. The notation is ID_ic, WBS_ic, SeedBS_ic and ID_th, WBS_th, SeedBS_th, respectively. In order to allow the competitors achieve their best performance, in some cases the inputs of the relevant functions as developed in R were adjusted. The competitors ID, WBS2, PELT and MARS were applied using the default values, while for DP_univar, the values proposed in the examples of the corresponding R function are used. For MOSUM, the function multiscale.localPrune was used, with the default values, which coincide with what is described in the simulation study of [16]. The maximum number of change-points to check for is required for the NOT algorithm and the number was chosen as $\max\{25,N+1\}$, where 25 is the default value. For MSCP, the value of the minimal window considered is chosen to be $\min\{\lceil\delta_{T}/2\rceil,20\}$, where 20 is the default value. The estimation of the standard deviation for CPOP is set to be done using the MAD method as described in Section 3. TF and SeedBS are employed based on the implementations found in https://github.com/hadley/l1tf and https://github.com/kovacssolt/SeedBinSeg/blob/master/SeedBS.R, respectively. Change-point detection using local_poly involves performing cross validation for every time series, for the optimal values to be chosen. The degree of the polynomial is set to 1, so that the fitted signals are piecewise-linear. Code on the simulations, as well as on how to implement DAIS, can be found in https://github.com/Sophia-Loizidou/DAIS. For the simulations we have used a modification of the DAIS algorithm described in Section 2, in order to make it more accurate. More specifically, DAIS restarts from the detected change-point instead of the start- and end-points of the interval where the detection occurred. This leads to an improvement of the accuracy of the algorithm without affecting the speed. This is also our recommendation on how to use the algorithm in practice.
@@
WBS [23] breakfast
WBS2 [25] breakfast
Piecewise-constant NOT [6] not
@@
local_poly [54] changepoints
As a measure of the accuracy of the estimated number of change-points, we present the difference between the number of change-points detected and the true number of change-points ($\hat{N}-N$). As a measure of accuracy of the location of the detected change-points, we give Monte-Carlo estimates of the mean squared error,
@@
where $T$ is the total length of the sequence and $\hat{f_{t}}$ is the ordinary least squares approximation of $f_{t}$ between two successive change-points. We also report the scaled Hausdorff distance, defined as
@@
where $n_{s}=\max_{j=1,\ldots,N}\tilde{\delta}_{j}$ is the length of the largest segment between successive change-points. The Hausdorff distance, $d_{H}$, is given for all signals except for signals that do not contain any change-points; in such cases, $d_{H}$ is uninformative. Finally, the average computational time for each method is also reported. The top performing algorithm, in terms of accuracy in the number of change-points detected, is presented in bold, along with all methods that performed within $5\%$. The signals used can be found in Appendix A and some further simulation results are in Appendix B.
@@
The signals included in this section contain more than one change-point. For signals with no or one change-point, please refer to Tables 8 and 10 in Appendix B. The signal (S1) (Table 3) is a particularly difficult structure for most algorithms because it includes a pair of consecutive change-points that shift the data sequence to opposite directions, which masks the existence of the change-points. Detection is more difficult when the change-points are close to each other and their isolation is unlikely or occurs only in a limited number of intervals, which is the case in this signal. The best performing method for (S1) is DAIS, having at least a $7.5\%$ improvement in the number of times the correct number of change-points is detected compared to all the other algorithms. The reason why DAIS performs very well is that it will most likely start the expansion around the change-points, as is discussed in Section 3.3, and so it has an advantage in detecting them, as the power of the contrast function $C_{s,e}^{b}(\boldsymbol{X})$ is maximized when $b$ is at the midpoint of $[s,e]$. For signals (S2) and (S3), the majority of the algorithms perform well; however, WBS_th and SeedBS_th tend to overestimate and MSCP seems to underestimate the number of the change-points, while DP_univar underestimates in signal (S2) and overestimates in signal (S3). For signals (S4), (S5) and (S6) in Tables 3 and 4, some methods struggle to detect the change-points, which are really close to each other. The signal (S4) includes change-points of different magnitudes and with different distances between them. ID_th and ID_ic perform well in different scenarios, but neither is consistent, which is expected as the thresholding approach is more appropriate for a large number of change-points, while the information criterion is better suited for a smaller number. Even though MOSUM performs well for signals with a small number of change-points, it fails to detect the change-points for (S4), (S5) and (S6). WBS2 performs well but is more computationally expensive compared to DAIS, with DAIS being 10-100 times faster than WBS2 in the simulations. Based on the simulations, we conclude that overall, DAIS performs very well in various different piecewise-constant signal scenarios, involving the number of change-points, the distance between them, as well as the magnitudes of the changes, when most competitors struggle to accurately detect the changes in some of the cases tested.
@@
Proceeding now with the piecewise-linear framework, the simulation results in Tables 5 and 6 show that in the case of piecewise-linear signals, DAIS performs at least as accurately as the algorithms ID_ic and CPOP, and better than other competitors available in the literature regarding the estimated number of change-points and their locations. The average computational time for DAIS is very low, and most of the times, significantly lower than the best-performing competitors. For signal (S10), DAIS outperforms all competitors except for CPOP in terms of the accuracy for the number of change-points detected. However, CPOP has a higher MSE and is at least four times slower compared to DAIS in all the examples presented.
@@
Table 3: Distribution of $\hat{N}-N$ over 100 simulated data sequences of the signals (S1), (S2), (S3) and (S4). The average MSE, $d_{H}$ and computational times are also given.
@@
Method Signal $\leq 3$ $-2$ $-1$ 0 1 2 $\geq 3$ MSE $d_{H}$ Time (s)
DAIS 0 15 0 81 2 2 0 0.014 0.089 0.004
ID_th 0 8 0 72 9 10 1 0.014 0.098 0.004
ID_ic 0 24 0 73 3 0 0 0.016 0.133 0.002
WBS_th 0 0 5 30 18 25 22 0.018 0.196 0.009
WBS_ic 0 30 0 68 2 0 0 0.016 0.166 0.025
WBS2 (S1) 0 10 2 75 7 1 5 0.015 0.091 0.446
PELT 0 77 0 23 0 0 0 0.026 0.397 0.001
NOT 0 26 0 72 2 0 0 0.016 0.146 0.024
MOSUM 0 15 0 74 6 4 1 0.014 0.108 0.006
MSCP 0 75 23 2 0 0 0 0.030 0.394 2.600
SeedBS_th 0 0 4 20 21 22 33 0.019 0.245 0.013
SeedBS_ic 0 32 0 66 2 0 0 0.017 0.176 0.013
DP_univar 0 0 0 6 3 10 81 0.045 0.334 0.138
@@
MSCP 100 0 0 0 0 0 0 0.949 0.182 0.072
SeedBS_th 0 0 1 66 27 6 0 0.025 0.018 0.012
SeedBS_ic 0 0 1 66 27 6 0 0.025 0.018 0.012
DP_univar 100 0 0 0 0 0 0 0.240 0.067 0.001
@@
MSCP 49 33 16 2 0 0 0 10.770 0.109 0.289
SeedBS_th 0 0 0 45 34 11 10 1.763 0.033 0.011
SeedBS_ic 0 0 0 96 4 0 0 1.599 0.010 0.011
DP_univar 0 0 0 0 0 0 100 4.544 0.112 0.004
DAIS 0 0 0 98 2 0 0 0.253 0.004 0.0014
ID_th 0 0 0 96 4 0 0 0.254 0.006 0.0008
ID_ic 22 11 42 24 1 0 0 1.540 0.111 0.0013
WBS_th 0 0 0 87 8 4 1 0.251 0.007 0.0009
WBS_ic 0 0 0 71 19 9 1 0.270 0.012 0.0018
WBS2 (S4) 0 0 0 99 1 0 0 0.235 0.004 0.0018
PELT 4 4 0 92 0 0 0 0.380 0.018 0.0001
NOT 0 0 0 92 7 1 0 0.242 0.007 0.0144
MOSUM 100 0 0 0 0 0 0 6.460 0.893 0.0001
MSCP 100 0 0 0 0 0 0 6.270 0.446 0.0257
SeedBS_th 0 0 3 86 6 5 0 0.267 0.008 0.0108
SeedBS_ic 0 0 0 73 20 5 2 0.265 0.011 0.0108
DP_univar 100 0 0 0 0 0 0 5.790 0.154 0.0001
Table 4: Distribution of $\hat{N}-N$ over 100 simulated data sequences of the signals (S5) and (S6). The average MSE, $d_{H}$ and computational times are also given.
@@
MSCP 0 100 0 3.930 0.057 1.400
SeedBS_th 0 0 100 0.454 0.009 0.012
SeedBS_ic 100 0 0 4.000 0.990 0.012
DP_univar 0 100 0 3.690 0.026 0.048
DAIS 0 0 100 0.234 0.003 0.013
ID_th 0 0 100 0.208 0.002 0.009
ID_ic 100 0 0 6.250 0.992 0.023
WBS_th 0 31 69 1.080 0.009 0.020
WBS_ic 100 0 0 6.250 0.992 0.020
WBS2 (S6) 0 0 100 0.199 0.001 0.284
PELT 19 38 43 1.670 0.082 0.001
NOT 100 0 0 6.250 0.992 0.041
MOSUM 100 0 0 6.250 0.992 0.003
MSCP 100 0 0 6.220 0.160 1.090
SeedBS_th 0 0 100 0.235 0.003 0.012
SeedBS_ic 100 0 0 6.250 0.992 0.013
DP_univar 100 0 0 5.710 0.025 0.030
Table 5: Distribution of $\hat{N}-N$ over 100 simulated data sequences of the signal (S8). The average MSE, $d_{H}$ and computational times are also given.
@@
Method Signal $-1$ 0 1 $2$ $\geq 3$ MSE $d_{H}$ Time (s)
DAIS 0 99 1 0 0 0.030 0.082 0.024
ID_th 0 71 28 1 0 0.036 0.130 0.024
ID_ic 0 95 5 0 0 0.031 0.096 0.008
CPOP 0 99 1 0 0 0.013 0.050 5.550
NOT (S8) 0 99 1 0 0 0.016 0.060 0.186
MARS 0 5 39 47 9 0.025 0.194 0.008
TF 0 0 0 0 100 0.018 0.440 0.768
TS 1 99 0 0 0 0.096 0.186 1.020
local_poly 0 1 5 12 82 0.054 0.503 125.400
Table 6: Distribution of $\hat{N}-N$ over 100 simulated data sequences of the signals (S9) and (S10). The average MSE, $d_{H}$ and computational times are also given.
@@
Method Signal $\leq-15$ $(-15,-2]$ $-1$ 0 1 $[2,15)$ $\geq 15$ MSE $d_{H}$ Time (s)
DAIS 0 0 0 100 0 0 0 0.266 0.295 0.026
ID_th 0 0 0 67 30 3 0 0.278 0.332 0.014
ID_ic 0 0 0 96 4 0 0 0.202 0.251 0.411
CPOP 0 0 0 91 9 0 0 0.381 0.247 0.439
NOT (S9) 100 0 0 0 0 0 0 4.730 99.00 1.440
MARS 100 0 0 0 0 0 0 4.700 98.50 0.003
TF 0 0 0 0 0 0 100 0.205 0.395 0.645
TS 0 33 37 30 0 0 0 0.854 0.580 1.430
local_poly 100 0 0 0 0 0 0 4.740 0.827 157.80
DAIS 0 0 0 100 0 0 0 0.038 0.196 0.016
ID_th 0 0 0 89 11 0 0 0.043 0.271 0.011
ID_ic 2 4 15 79 0 0 0 0.117 2.860 0.365
CPOP 0 0 0 99 1 0 0 0.252 0.291 0.062
NOT (S10) 100 0 0 0 0 0 0 1.060 119.000 0.050
MARS 100 0 0 0 0 0 0 1.060 118.000 0.002
TF 0 0 0 0 0 1 99 0.191 0.307 0.300
TS 100 0 0 0 0 0 0 1.060 119.000 0.480
local_poly 100 0 0 0 0 0 0 1.060 76.700 16.900
6 DAIS extensions
6.1 Temporal dependence
We first consider relaxing the assumption of independence for the error terms $\epsilon_{t}$. Temporal dependence is therefore introduced in the observed data sequence. In such cases, methods that reduce the autocorrelation in the time series can be applied. This can be done using a subsampling technique, which is briefly described below.
For a chosen integer $s$ we subsample the observed time series $X_{t}$ by choosing every $s^{\text{th}}$ observation. This creates $s$ different data sequences, for which the autocorrelation is reduced compared to the original data sequence. If, for example, $s=5$ and $T=1000$, the new data sequences are of the form $\{X_{1},X_{6},\ldots,X_{996}\}$, $\{X_{2},X_{7},\ldots,X_{997}\}$, $\{X_{3},X_{8},\ldots,X_{998}\}$, $\{X_{4},X_{9},\ldots,X_{999}\}$, $\{X_{5},X_{10},\ldots,X_{1000}\}$. DAIS is applied to each one of these smaller data sequences and we obtain $s$ sets of detected change-points. We can then apply a majority voting rule, which discards any change-points that are not detected in at least $\eta$ time series, where $\eta\leq s$ is a positive integer. The change-points are then transformed to represent locations in the original time series. The purpose of taking every $s^{\text{th}}$ observation is to not include the observations which have the highest correlation between them. That is, the observations just before or just after each data point. Thus, for larger values of $s$, we manage to reduce the autocorrelation in the resulting disjoint data sequences, but we obtain a larger number of smaller data sequences, which has a negative impact on the detection accuracy. More details on this can be found in [1]. A different way to reduce autocorrelation is using a pre-averaging technique, such as the one described in Section 6.2, which involves averaging the data sequence before applying the algorithm in order to reduce the correlation between consecutive observations.
6.2 Heavy-tailed noise
In this subsection we relax the assumption that the error term $\epsilon_{t}$ is Gaussian and we instead consider a heavy-tailed distribution for $\epsilon_{t}$. To improve the performance of DAIS in such cases, the data sequence can be pre-processed by averaging the time series to bring the noise closer to Gaussianity, as described in Section 4.5 of [2]. Using this technique, we take advantage of the Central Limit Theorem, the noise of the observations with large $\epsilon_{t}$ is reduced in absolute value, and we obtain a smaller data sequence with less extreme values for all the data points. The idea is that for a chosen integer $s$, we define $Q=\lceil T/s\rceil$ and set
$X^{\ast}_{q}=\frac{1}{s}\sum_{t=(q-1)s+1}^{qs}X_{t},\text{ for }t=1,2,\ldots,q\quad\text{and}\quad X^{\ast}_{Q}=\frac{1}{T-(Q-1)s}\sum_{t=(Q-1)s+1}^{T}X_{t}.$
DAIS can then be applied on $X^{\ast}_{t}$ to obtain change-points $\hat{r}^{\ast}_{1},\ldots,\hat{r}^{\ast}_{\hat{N}}$. The change-points of the original time series $X_{t}$ can be obtained using the transformation $\hat{r}_{i}=(\hat{r}^{\ast}_{i}-1)s+\lfloor\frac{s+1}{2}\rfloor$, for $i=1,\ldots,\hat{N}$. The choice of the value of $s$ is important. On the one hand, taking larger values brings the data closer to Gaussianity. On the other hand, these values of $s$ return smaller data sequences that have been pre-processed more and so we lose on the accuracy of both the location and the number of change-points detected. In practice we recommend $s=5$.
The performance of DAIS in the case of heavy-tailed data can be found in Table 7. For the pre-averaging technique we use $s=5$ and the heavy-tailed noise is chosen to be the Student’s t-distribution with degrees of freedom being equal to either 5 or 7. Signals (S1) and (S7) are piecewise-constant and (S8) is piecewise-linear and their description can be found in Appendix A. As expected, the accuracy of DAIS is reduced compared to Section 5. Moreover, the more heavy-tailed the noise, the more the overdetection. This is again expected, as the spikes caused by the heavy-tailed distribution of $\epsilon_{t}$ can be mistakenly detected as change-points by the algorithm, even after the pre-processing. However, DAIS performs relatively well for signal (S1) even for 5 degrees of freedom.
Table 7: Distribution of $\hat{N}-N$ over 100 simulated data sequences of the signals (S1), (S7) and (S8) with noise following the Student’s t-distribution with $d=5,7$ degrees of freedom. The average MSE, $d_{H}$ and computational times including the pre-averaging are also given.